OpenStudy (anonymous):

Locate all critical points of 6x^2-2x+1 can you please explain how as well? plz and thank you

OpenStudy (cwrw238):

have you done any calculus?

OpenStudy (anonymous):

yes, I recently got to this point in time of class, and I forgot how to do it, and Im a bad note taker :/

OpenStudy (cwrw238):

ok you first find the derivative of the function which gives your the slope of tangent at any point on the graph. At a turning point this slope = 0. f(x) = 6x^2 - 2x + 1 can you find f'(x) ?

OpenStudy (anonymous):

f'(x)=12x-2

OpenStudy (cwrw238):

right so 12x = 2 x = 1/6 other coordinate y = 6(1/6) - 2(1/6) + 1 = 5/3 so theres a turning point at(1/6, 5/3) do they want to know its nature ( maximum, minimum os point of inflection)?

OpenStudy (anonymous):

they dont want to, right now, but possibly in the future, if you can explan thatll be awesome

OpenStudy (cwrw238):

ok there's 2 ways either find the slope before and after the critical point or check sign of the second derivative

OpenStudy (cwrw238):